P(A speaks truth is) = 4/5. A coin is tossed. A reports that a head appears. The probability that actually there was a head is.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
By applying bayes theorem
probability of head or tail occuring=1/2
p(A speaks truth when head occurs)=\frac{4/5(1/2)}{(4/5)(1/2)+(1/5)(1/2)}
=4/5